OpenStudy (anonymous):

Mrs.Tetic bought 4 3/4 yards of fabric she needs another 4 3/8. how many yards of fabric does she need all together? 4 3/4, 9 1/8, 4 3/8, 8 3/4, I think its 9 and 1/8! Im not sure though XD

OpenStudy (anonymous):

=3 (3/4) + 4 (3/8) =(15/4) + (35/8) =(30+35)/8 =65/8 =8 (1/8)

OpenStudy (anonymous):

Hmmm I have nine and 1/8 not 8 and 1/8.. but could the answers they gave be wrong?

OpenStudy (anonymous):

oh wait, i had a typo. it's suppose to be nice 1/8. sorry

OpenStudy (anonymous):

Its ok!